OpenStudy (anonymous):

Eliminate the parameter. x = t2 + 2, y = t2 - 4

OpenStudy (agent0smith):

\[\Large x = t^2 + 2\]can you solve this for t?

OpenStudy (anonymous):

t^2 = x - 2 y = (x-2) - 4 = x - 6

OpenStudy (agent0smith):

Actually solve for t^2, not t. Since you need t^2 in y = t^2 - 4 anyway

OpenStudy (anonymous):

wait confused

OpenStudy (anonymous):

x-2=t^2

OpenStudy (agent0smith):

Correct. Now replace the t^2 in y = t^2 -2 with that.

OpenStudy (anonymous):

y=(x-2)-2

OpenStudy (agent0smith):

Good, now just simplify it.

OpenStudy (anonymous):

poop, help lol

OpenStudy (agent0smith):

lol you know how to simplify this :P

OpenStudy (anonymous):

im so confused because im n ot getting any of the answers

OpenStudy (anonymous):

y = x - 6, x ≥ 1 y = x + 6, x ≥ 1 y = x2 - 6, x ≥ 1 y = x2 + 6, x ≥ 1

OpenStudy (anonymous):

like does are the answers

OpenStudy (agent0smith):

x = t2 + 2, y = t2 - 4 you got t^2 = x-2 correct, but used... y=(x-2)-2 It's not y = t^2 - 2 ;)

OpenStudy (anonymous):

so x-2-2

OpenStudy (agent0smith):

No no, it's not y = t^2 - 2 (which you used) it's y= t^2 - 4

OpenStudy (anonymous):

ohhhh

OpenStudy (agent0smith):

Got it? :D

OpenStudy (anonymous):

nope lol i got -4x+8

OpenStudy (agent0smith):

lol wait. y= t^2 - 4 and t^2 = x-2

OpenStudy (anonymous):

yeah no not getting this

OpenStudy (agent0smith):

Okay. You know y = t^2 - 4 and t^2 = x-2 right? So replace t^2 in y = t^2 - 4, with x-2.

OpenStudy (anonymous):

okay so y=(x-2)-4

OpenStudy (agent0smith):

Yes :) you had it before... you just used the wrong equation :)

OpenStudy (anonymous):

okay now from there what?

OpenStudy (agent0smith):

I know you can simplify this: y=(x-2)-4 (there's no need for the brackets now)

OpenStudy (anonymous):

x-6

OpenStudy (agent0smith):

Yep

OpenStudy (anonymous):

yay!

OpenStudy (agent0smith):

:D yay!

OpenStudy (anonymous):

Eliminate the parameter. x = 3t, y = t + 7 would this be sorta the same

OpenStudy (agent0smith):

Yes, solve for t in terms of x first, then plug it into the y equation

OpenStudy (anonymous):

y = x /3+ 7?

OpenStudy (agent0smith):

Yes :D

OpenStudy (anonymous):

AHH YOU ARE TO AMAZING HAAH

OpenStudy (agent0smith):

haha well you learn fast, good job!

OpenStudy (anonymous):

Eliminate the parameter. x = 4 cos t, y = 4 sin t

OpenStudy (anonymous):

x^2 = 16 cos^2 t y^2 = 16 sin^2 t x^2 + y^2 = 16 cos^2 t + 16 sin^2 t x^2 + y^2 = 16 (cos^2 t + sin^2 t) x^2 + y^2 = 16(1) x^2 + y^2 = 16

OpenStudy (agent0smith):

^ is that your working?

OpenStudy (anonymous):

Someone was explaining it to me

OpenStudy (agent0smith):

Do you follow?

OpenStudy (anonymous):

What do you mean?

OpenStudy (agent0smith):

Understand it?

OpenStudy (anonymous):

no :( Like honestly you've been the only person that has made me get math

OpenStudy (agent0smith):

This one is a bit different to before... you can't just solve for one now. x = 4 cos t, y = 4 sin t Square both equations (just square everything in both of them) in the first step, since it'll come in handy later x^2 = 16 cos^2 t y^2 = 16 sin^2 t

OpenStudy (anonymous):

okay got that

OpenStudy (agent0smith):

Now add them together, so add everything on the left side and the right side x^2 = 16 cos^2 t y^2 = 16 sin^2 t to get x^2 + y^2 = 16cos^2 t + 16 sin^2 t

OpenStudy (anonymous):

factor now?

OpenStudy (agent0smith):

Yes, factor out the 16 on the right... and remember sin^2 t + cos^2 t = 1

OpenStudy (anonymous):

so it would just be 16(1) yay i got it right

OpenStudy (anonymous):

then just x^2 + y^2 = 16

OpenStudy (agent0smith):

yes! :D

OpenStudy (anonymous):

♥♥ thanks

OpenStudy (agent0smith):

♥♥ welcome